Hot water at a temperature of t is added to twice that amount at a temperature of 30degree centigrade, if the resulting temperature of the mixture is 50degree centigrade , calculate t
1
Expert's answer
2020-12-06T17:27:50-0500
"mc(T-50)=2mc(50-30)\\T=50+2(50-30)=90\degree C"